You are ready to buy a house, and you have $20,000 for a down payment and closing costs. Closing costs are estimated to be 4% of the loan value. You have an annual salary of $36,000, and the bank is willing to allow your monthly mortgage payment to be equal to 28% of your monthly income. The interest rate on the loan is 6% per year with monthly compounding (.5% per month) for a 30-year fixed rate loan. How much money will the bank loan you? How much can you offer for the house?
Answer #1

Monthly payment = annual salary/12*0.28

=36000/12*0.28 = 840

PVOrdinary Annuity = C*[(1-(1+i/(f*100))^(-n*f))/(i/(f*100))]
C = Cash flow per period
i = interest rate
n = number of payments I f = frequency of payment
PV= 840*((1-(1+ 6/1200)^(-30*12))/(6/1200))
PV = 140104.96
Using Calculator: press buttons "2ND"+"FV" then assign
PMT =840
I/Y =6/12
N =30*12
FV = 0
CPT PV
Using Excel
=PV(rate,nper,pmt,FV,type)
=PV(6/(12*100),12*30,,PV,)